# The Ultimate Configuration Management Tool? Lessons from a Mixed Methods Study of Ansible’s Challenges

### v2

@@ 10. Conclusion
We combined a large-scale quantitative analysis of 59,157 online Ansible discussions with 20 practitioner interviews to identify key challenges, contextualize their real-world impact, and propose solutions. We found issues related to control flow, performance bottlenecks, limited debugging capabilities, and inconsistent documentation, highlighting the tensions among usability, flexibility, and scalability within Ansible’s ecosystem. We offer concrete recommendations and show how Ansible can better evolve to serve novice users and large-scale deployments. These contributions, alongside broader reflections on community-driven knowledge, provide a replicable roadmap for assessing and improving IaC solutions.

### v3

@@ 10. Conclusion
We combined a large-scale quantitative analysis of 59,157 online Ansible discussions with 20 practitioner interviews to identify key challenges, contextualize their real-world impact, and propose solutions. We found issues related to control flow, performance bottlenecks, limited debugging capabilities, and inconsistent documentation, highlighting the tensions among usability, flexibility, and scalability in Ansible’s ecosystem. We offer concrete recommendations and show how Ansible can better serve users and large-scale deployments. These contributions, alongside broader reflections on community-driven knowledge, provide a replicable roadmap for assessing and improving IaC solutions.